OpenStudy (anonymous):

find the derivative of- (x+s)(sqrt(3s^2+2sx-x^2))/4

OpenStudy (anonymous):

\[((x+s)(\sqrt{3s ^{2}+2sx-x ^{2}}))\div2\]

OpenStudy (mathstudent55):

Is it/4 or /2?

OpenStudy (anonymous):

derivative with respect to what? d/dx or d/ds

OpenStudy (anonymous):

/4 sorry

OpenStudy (anonymous):

and d/dx

OpenStudy (anonymous):

everytime I attempt it, it gets real messy, real fast

OpenStudy (mathstudent55):

\(u = x + s\) \( v = (s^2 + 2sx - x^2)^{ \frac{1}{2} } \) \(f(x) = uv\) \(f'(x) = uv' + vu' \)

OpenStudy (mathstudent55):

\(u' = 1\) \(v' = \dfrac{1}{2}(s^2 + 2sx - x^2)^{-\frac{1}{2}} (2s - 2x) = (s^2 + 2sx - x^2)^{-\frac{1}{2}} (s - x) \)

OpenStudy (anonymous):

there was a three in front of the s^2

OpenStudy (anonymous):

ohhhh I got it from here

OpenStudy (anonymous):

thanks a bunch

OpenStudy (mathstudent55):

Wlcm. Remember the 3 I left out, and remember the /4 which is just a constant.